OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of . y - 5 = (x + 1) y + 1 = (x - 5) y + 5 = -(x - 1) y - 1 = (x + 5)

OpenStudy (noseboy908):

all yo need to know is that slope-intercept form is "y-y1=m(x-x1)

OpenStudy (noseboy908):

*you

OpenStudy (anonymous):

the slope is \[\frac{ 3 }{ 2 }\]

OpenStudy (noseboy908):

is that what's given?

OpenStudy (anonymous):

yess

OpenStudy (noseboy908):

in that case, none of your possible given answers are correct.